OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

What are you having trouble with?

OpenStudy (anonymous):

It means that the equation = 0 if you plugin that number. e.g. Solution to 2x-10 is 5, since 2(5)-10 =0 Solutions to (x-3)(x+4) are 3, and -4, since (3-3)(3+4)=0 and (-4-3)(-4+4)=0 and so on...

OpenStudy (anonymous):

2x-10 and (x-3)(x+4) are expressions, not equations. Your statement would be true for 2x-10=0 and (x-3)(x+4)=0, which is fundamentally different. A solution to the equation demonstrates equality in the equation.

OpenStudy (anonymous):

Ok thanks. Can u also help me with my problem?

OpenStudy (anonymous):

Solutions make statements true.